OpenStudy (anonymous):

find the tangent of the graph f(x,y)=6x-3x^2-y^2 at th point (1,2,-1).how ?pls help

OpenStudy (amistre64):

youll need to take the derivative; and is that a tangent plane? or tangent line?

OpenStudy (anonymous):

tANGENT PLANE

OpenStudy (amistre64):

find the gradient sounds familiar then

OpenStudy (amistre64):

the gradient produces a normal vector at a given point ...

OpenStudy (anonymous):

but how i can take derivatives of z co ordinate?

OpenStudy (amistre64):

z = f(x,y)=6x-3x^2-y^2 f(x,y,z)=6x-3x^2-y^2-z now the partials form the gradient; (Fx, Fy, Fz)

OpenStudy (amistre64):

Fx = 6-6x Fy = -2y Fz = -1 at the point (1,2,-1) we get Fx = 0 Fy = -4 Fz = -1

OpenStudy (anonymous):

i can do the rest but the prblem is how does f(x,y,z)=the equation happen?

OpenStudy (amistre64):

recall how y = f(x) earlier in the courses?

OpenStudy (amistre64):

the usual notation in 3d space is: z is a function of x and y, z = f(x,y)

OpenStudy (anonymous):

ok thnx a lot

OpenStudy (amistre64):

youre welcome